A spark will jump between two people if the electric field exceeds 4.0 x 106 V/m. You shuffle across a rug and a spark jumps when you put your finger 0.48 cm from another person’s arm. Calculate the potential difference between your body and the other person’s arm.
1
Expert's answer
2020-03-02T09:07:41-0500

Solution. The relationship of the electric field with the potential difference is related by the formula


"\Delta φ= Ed"

where ∆φ is potential difference; E= 4.0 x 106 V/m is electric field; d=0.48 cm is distance. Therefore


"\Delta φ= 4.0\times 10^6\times 0.48\times 10^{-2}=19200V"

Answer.


"\Delta φ=19200V"